OpenStudy (anonymous):

i am trying to reduce this further… what do i do.. y^'= 2 sinx(cos(x)sin(2x)+sin(x)cos(2x))

OpenStudy (anonymous):

\[y \prime= 2sinx(cosxsin2x+sinxcos2x)\]

OpenStudy (btaylor):

The part with cos(x)sin(2x)+sin(x)cos(2x) looks like the angle addition identity for sine: \[\sin(\alpha + \beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\] So, you can simplify that chunk to sin(x+2x), or sin(3x).

OpenStudy (anonymous):

my final answer should be \[2\sin^2(x)(2\cos(2x)+1)\]

OpenStudy (btaylor):

where do you get the sin^2?

OpenStudy (btaylor):

I'd get: \[2 \sin (x) \sin(3x)\]

OpenStudy (anonymous):

no idea. he gives us answers to the problem but we have to show steps and that is answer he has given

OpenStudy (btaylor):

I'm not sure how to get there...sorry!

OpenStudy (anonymous):

thanks anyways! i don't think he uses identities.